OpenStudy (anonymous):

Convert the following x+5y=-3 from standard form to slope intercept form

Nnesha (nnesha):

slope intercept form is y = mx + b where m is slope which is always with x and b is y-intercept so you have to solve for y

Nnesha (nnesha):

first subtract x both side let m know what you get after this :)

Nnesha (nnesha):

me**

Nnesha (nnesha):

 what you get ??

OpenStudy (alexandervonhumboldt2):

frst subtratc x from both sdes

OpenStudy (alexandervonhumboldt2):

5y=-3-x

OpenStudy (alexandervonhumboldt2):

now dvideit by 5 y=(-3-x)/5 convert: y=-x/5-3/5

OpenStudy (anonymous):

I'm sorry I'm having a hard time seeing the equation i only see half

OpenStudy (alexandervonhumboldt2):

hmm relaoad your page

OpenStudy (alexandervonhumboldt2):

y=(-3-x)/5 y=-x/5-3/5

Nnesha (nnesha):

@flirty45 i'm waiting 4 ur reply let me know what you get :)

OpenStudy (anonymous):

5y=-3-x

Nnesha (nnesha):

yes right now divide both side by 5

OpenStudy (anonymous):

x=-3/5-5

Nnesha (nnesha):

oops who kidnapped our x :D??????

OpenStudy (anonymous):

x=3/5-x

Nnesha (nnesha):

 both side dividing by 5 and remember 3 is negative on right side